OpenStudy (anonymous):

Radical Equations: What should you multiply the entire equation by when there are fractions?

OpenStudy (anonymous):

what the equation?

OpenStudy (johnweldon1993):

Give me an example If all the fractions have the same denominator you would multiply everything by that because \[\large 5\times (\frac{1}{5} + \frac{2}{5})\]will become \[\large \cancel{5}\times (\frac{1}{\cancel{5}} + \frac{2}{\cancel{5}})\] which is just \[\large 1 + 2\]

OpenStudy (anonymous):

10/30=20/60

OpenStudy (johnweldon1993):

\[\large \frac{10}{30} = \frac{20}{60}\] Well for that you would just, evaluate it....or cross multiply to see its true, no need to multiply both sides by something

OpenStudy (anonymous):

10/30 = 20/60 - due to the numerators and denominators equaling 20/60 when multiplied by 2/2 both sides of the proportion are similar